OpenStudy (anonymous):

Use a half–angle identity to find the exact value of sin 105°.

OpenStudy (anonymous):

105=60+?

OpenStudy (anonymous):

105=60+45

OpenStudy (anonymous):

now we use \[\sin (A+B)=\sin Acos B+\sin Bcos A\]

OpenStudy (anonymous):

wait the question says half angle this is compound angle

OpenStudy (anonymous):

sin(60+45) = sin60cos45 + sin60cos45

OpenStudy (anonymous):

this way will work but the question is looking for the 1/2 angle identity

OpenStudy (anonymous):

oh my gosh! YES! hahaha tats what i cant figure out the half angle identity to use for this problem.

OpenStudy (anonymous):

*thats

OpenStudy (anonymous):

\[\sin(A)=\sqrt{\frac{ 1-\cos 2A }{ 2 }}\]

OpenStudy (anonymous):

so let A=105 then 2A=210

OpenStudy (anonymous):

so it would be -1/2 radical 2+radical 3 ?

OpenStudy (anonymous):

2+radical3 being under a radical

OpenStudy (anonymous):

if you have a calculator find cos(210) or you rewrite it as cos(180+30)="-(sqrt(3)/2)". So plugging this value into jonask's equation will give you the answer

OpenStudy (anonymous):

1/2 2+radical 3 . thats my answer

OpenStudy (anonymous):

this one is right :) \[\sqrt{2+\sqrt{3}}\div2 \]

OpenStudy (anonymous):

so i was close ! :) thank you !

OpenStudy (anonymous):

your welcome.. gud bye..